Solve the system of linear equations using addition/subtraction method.
2x+y=3
x-2y=-1

Answer by padejr(3)

2x+y = 3
x-2y = -1
use elemination method, we multiply equation 1 by 2
we get 4x+2Y = 6
Add equation 2 and the last one
we get 5x = 5
x = 1
Substitute x in to either equation 1 or 2
we get y = 1